OpenStudy (anonymous):

Simplify. Assume all variables represent positive real numbers. sqrt16x^2

OpenStudy (anonymous):

\[\sqrt{16x²} = \sqrt{4²x²} = 4x\]

OpenStudy (anonymous):

sqrt(16x^2) = sqrt(16)sqrt(x^2) = 4sqrt(x^2) since x > 0 , 4sqrt(x^2) = 4x